# Towards Cross-Lingual LLM Evaluation for European Languages

### v1

@@ Appendix B European LLM Evaluation Results
Model Name Avg. ARC HellaSwag
@@
Mistral-Nemo-Base-12.2B_2407202020https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .562 .615 ± .071 .641 ± .094
Mistral-Nemo-Instruct-12.2B_2407212121https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407 .598 .621 ± .071 .624 ± .093
@@
Occiglot-7b-eu5222222https://huggingface.co/occiglot/occiglot-7b-eu5 .404 .470 ± .130 .511 ± .143
Occiglot-7b-eu5-Instruct [^23] https://huggingface.co/occiglot/occiglot-7b-eu5-instruct .406 .484 ± .132 .519 ± .146
Pharia-1-LLM-7B-control  [^24] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control .333 .393 ± .174 .433 ± .181
Pharia-1-LLM-7B-control-aligned [^25] https://huggingface.co/Aleph-Alpha/Pharia-1-LLM-7B-control-aligned .346 .396 ± .177 .438 ± .185
@@
Model Name MMLU TruthfulQA GSM8k
@@
Model Name Avg. ARC HellaSwag
@@
Model Name MMLU TruthfulQA GSM8k
@@
Model Name Avg. ARC HellaSwag
@@
Model Name MMLU TruthfulQA GSM8k
@@
Model Name Avg. ARC HellaSwag
@@
Model Name MMLU TruthfulQA GSM8k
@@
c4ai-command-r-35B-v01 .540 ± .045 .542 ± .021 .424 ± .053
Model ARC GSM8K HellaSwag MMLU TruthfulQA
@@
c4ai-command-r-35B-v01 .201 .124 .205 .112 .030
Table 16: Differences between high and medium resource languages
Model Name PC
Aya-23-8B 0.571
Meta-Llama-3-8B-Instruct 0.562
Mistral-7B-Instruct-v0.3 0.562
Meta-Llama-3-8B 0.556
gemma-7b 0.555
Llama-2-7b-chat-hf 0.553
Meta-Llama-3.1-8B 0.552
gemma-1.1-7b-it 0.552
Meta-Llama-3.1-8B-Instruct 0.549
occiglot-7b-eu5 0.548
occiglot-7b-eu5-instruct 0.546
Mistral-7B-v0.1 0.542
Mistral-7B-Instruct-v0.2 0.539
Mistral-7B-v0.3 0.536
Phi-3-mini-4k-instruct 0.534
Mistral-7B-Instruct-v0.1 0.528
Llama-2-7b-hf 0.520
bloomz-7b1 0.439
bloom-7b1 0.401
Table 17: Model Performance Correlation Matrix (Sorted by Correlation Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
Language PC
es 0.589991
fr 0.586389
it 0.567793
sv 0.564892
da 0.557272
nl 0.555682
ro 0.546222
de 0.545930
sk 0.473632
hu 0.444939
Table 18: Language Correlation Coefficients (Sorted by Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
Task PC
EU21-ARC 0.610
EU21-HeSw 0.543
EU21-MMLU 0.538
Table 19: Task Correlation Coefficients (Sorted by Coefficient in Descending Order)
(PC: Pearson Correleation Coefficient)
@@
Appendix C Correlation with Human Preference Scores
@@
Model and Hugging Face Link
c4ai-command-r-35b-v01
gemma-1.1-7b-it
gemma-2-27b-it
gemma-2-9b-Instruct
Llama-2-7b-chat-hf
Llama-2-13b-chat-hf
Meta-Llama-3-8B-Instruct
Meta-Llama-3.1-8B-Instruct
Meta-Llama-3.1-70B-Instruct
Mistral-7B-Instruct-v0.1
Mistral-7B-Instruct-v0.2
Mixtral-8x7B-Instruct-v0.1
Phi-3-medium-14b-4k-instruct
Phi-3-mini-128k-Instruct
Phi-3-mini-4k-Instruct
Vicuna-13b-v1.5
Vicuna-33b-v1.3
Table 21: Model Overview
@@ Appendix D Influence of translation service
Table 29: Comparison of ogx and okapi scores for ARC and Hellaswag benchmarks (Higher values in bold)
Language ARC Hellaswag MMLU
OGX OKAPI OGX OKAPI OGX OKAPI
@@
Table 30: Average sample number for each task of eleven translations.
Split Original OGX OKAPI
MMLU Val. 1531 276.18 ±$\pm 8$
Test 14042 13198.09 ±$\pm 101$
Hellaswag Val 10042 10035.18 ±$\pm 19$ 9276.63 ±$\pm 112$
ARC Val. 299 299 272.66 ±$\pm 0.5$
Test 1172 1172 1071 ±$\pm 2$
TruthfulQA Val. 817 817 782.09 ±$\pm 6$
Table 31: Hellaswag translation examples compared between OGX and OKAPI
@@
German/OGX COMET: 0.6329 Wie man einen guten Wellensittich auswählt. Suchen Sie nach Wellensittichen, die von einem seriösen Verkäufer verkauft werden. Sie können Wellensittiche im Internet, in Kleinanzeigen oder in Zoohandlungen finden, die zum Verkauf stehen. Sie können auch einen Wellensittich aus Ihrem örtlichen Tierheim bekommen. - Die meisten Tierheime verkaufen Wellensittiche gegen eine Pauschalgebühr, so dass sich Ihre Bankkosten in Grenzen halten. Vielleicht können Sie auch die Preise vergleichen, um ein gutes Angebot zu bekommen. - Sie müssen einen angemessenen Preis finden. Im Allgemeinen sind 20-50 % des Schätzpreises angemessen. - Wo auch immer Sie Ihren Wellensittich kaufen, stellen Sie sicher, dass der Verkäufer gesunde und artgerecht gehaltene Vögel hat. Schauen Sie sich die Online-Bewertungen der Verkäufer an, die Sie finden. - Grundsätzlich müssen Sie einen guten Verkäufer finden, der bereits Vorbesitzer hatte. Erkundigen Sie sich, ob der Wellensittich den Eltern gehört hat.
@@
French/OGX COMET: 0.5770 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera celle d’un joueur de saxophone ténor, ainsi que le site web permettant de recevoir des leçons. Un homme noir apparaît ensuite, avec un pull vert et un chapeau noir, jouant du saxophone. l’homme - bluffe à la batterie et la vidéo continue, à différents moments, à montrer des personnes jouant de l’instrument. - - continue à jouer et applaudit, serre la main de l’homme, lui lance une serviette, écrit sur la serviette et lui donne une carte. pose ensuite l’instrument sur ses genoux et commence à parler de ce qu’il fait. - saute dans une vidéo en noir et blanc et s’arrête temporairement de jouer en disant " big calibers " (gros calibres).
@@
Spanish/OGX COMET: 0.7803 Un hombre bebe de una lata junto a un agujero en la nieve. él - nada lentamente por una masa de agua poco profunda. - habla y gesticula a la cámara. - un joven le muerde los dientes. - se acerca con pico y pala.
@@
Italian/OGX COMET: 0.8592 Come alleviare un virus intestinale. Evitate i cibi solidi per qualche ora per permettere allo stomaco di calmarsi. Quando si avvertono i primi sintomi di un virus intestinale, è necessario lasciare che lo stomaco si stabilizzi. Mangiare cibo subito potrebbe irritare maggiormente la mucosa gastrica, peggiorando i sintomi. - Per qualche ora, evitate i cibi solidi. Bevete piccoli sorsi d’acqua o masticate del ghiaccio. - Evitare di mangiare qualcosa per almeno 10 minuti. Gli studi non hanno riscontrato che mangiare per qualche ora possa alleviare i sintomi dello stomaco. - Cercare di mangiare un pasto solido qualche ora prima dell’attacco di nausea non farà altro che peggiorare i sintomi. Se vi sentite male dopo l’inizio dei sintomi, smettete di mangiare. - Cercate di mangiare qualche ora prima del lavoro o della scuola, soprattutto se vi sentite male. Anche muoversi un po’ può ridurre la nausea e aiutare lo stomaco a stabilizzarsi.
@@
Swedish/OGX COMET: 0.8174 En man med hörlurar befinner sig i en studio och talar i en mikrofon. Andra män visas, varav en börjar spela munspel. han - slutar spela och säger något som får dj:n att skratta. - ses leka intensivt med kameran. - leker med en jack o lantern och lämnar rummet. - verkar mycket nöjd med den information han har samlat in när han låter programledaren tala.
@@
Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
@@
Slovak/OGX COMET: 0.8336 Muž beží a pritom drží tyč na dráhe. on - používa tyč na skok vo vzduchu. - beží po dráhe a obopína nohy okolo tyče. - vbehne do pieskoviska a začne skákať. - dopadne na hromadu piesku.
@@
Danish/OGX COMET: 0.8450 En lille pige løber væk fra pinataen, mens hun holder pinden. Barnet vender tilbage og rammer pinataen én gang. personerne - fortsætter med at fjerne nøgleingredienser fra poserne. - jubler over barnet, der bliver ramt. - ser til, mens pigen kaster pinden to gange. - alle klapper, og pigen smiler og vender sig mod kameraet.
@@
Hungarian/OGX COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
@@
Table 32: one random ARC and MMLU translation examples each compared between OGX and OKAPI
@@
German/OGX COMET: 0.8345 Auf wen trafen die anatomisch modernen Menschen bei ihrer ersten Ankunft im Nahen Osten? - Neandertaler, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo erectus, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo sapiens, der sich unabhängig von den Menschen in Asien entwickelt hatte und dort vor 200.000 Jahren ankam - Alle oben genannten
@@
German/OGX COMET: 0.8546 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält. Welcher dieser Schritte sollte zuerst erfolgen? - Wiederhole die Untersuchung. - Schreibe einen Bericht über die Ergebnisse. - Fertige eine Tabelle an, um die Daten festzuhalten. - Beobachte täglich die Glühbirnen.

### v2

@@ B.3 Downstream results tables
Model Name Avg. EU21-ARC EU21-HeSw
@@
Mistral-Nemo-Base-12.2B_2407 .562 .615 ± .071 .641 ± .094
Mistral-Nemo-Instruct-12.2B_2407 .598 .621 ± .071 .624 ± .093
@@
Occiglot-7b-eu5 .404 .470 ± .130 .511 ± .143
Occiglot-7b-eu5-Instruct .406 .484 ± .132 .519 ± .146
Pharia-1-LLM-7B-control .333 .393 ± .174 .433 ± .181
Pharia-1-LLM-7B-control-aligned .346 .396 ± .177 .438 ± .185
@@
Model Name EU21-MMLU EU21-TQA EU21-GSM8k
@@
Model Name Avg. EU21-ARC EU21-HeSw
@@
Model Name EU21-MMLU EU21-TQA EU21-GSM8k
@@
Model Name Avg. EU21-ARC EU21-HeSw
@@
Model Name EU21-MMLU EU21-TQA EU21-GSM8k
@@
Model Name Avg. EU21-ARC EU21-HeSw
@@
Model Name EU21-MMLU EU21-TQA EU21-GSM8k
@@
Table 20: Differences between high and medium resource languages
EU21 EU21- EU21- EU21- EU21-
Model ARC GSM8k HeSw MMLU TQA
@@
B.4 Downstream results heatmaps
@@
Appendix C Correlation with Human Preferences
C.1 Battle dataset overview
The battle dataset [^20] https://storage.googleapis.com/arena_external_data/public/clean_battle_20240814_public.json (as of 2024-10-15) dataset we used for computing the Elo scores was published by LMSYS together with a Jupyter notebook [^21] https://colab.research.google.com/drive/1KdwokPjirkTmpO_P1WByFNFiqxWQquwH (as of 2024-10-15)for reproducibility purposes. Table 21 shows the proportions of the different languages in the dataset, according to the classification contained in the dataset itself:
@@ C.1 Battle dataset overview
C.2 Computation of model rank
The model rank for a model $m_{A}$ is computed by the FastChat codebase as 1 plus the number of models $m_{B}$ such that the 97.5% quantile of the Elo score of $m_{A}$ is lower than the 2.5% quantile of the Elo score of $m_{B}$.
For each such $m_{B}$, there is a 97.5% probability that the “true” Elo score of $m_{B}$ is greater than its 2.5% quantile, and this quantile is greater than the 97.5% quantile of the Elo score of $m_{A}$, which in turn has a 97.5% probability of being greater than the “true” Elo score of $m_{A}$. Thus, there is a probability greater than $97.5\%*97.5\%=95.0625\%$ that the “true” Elo score of $m_{B}$ is greater than the “true” Elo score of $m_{A}$.
C.3 Elo Score Tables
Tables LABEL:tab:lmsys_pearson_arc_few_shot, LABEL:tab:lmsys_pearson_gsm8k_few_shot, LABEL:tab:lmsys_pearson_hellaswag_few_shot, LABEL:tab:lmsys_pearson_mmlu_few_shot, and LABEL:tab:lmsys_pearson_truthfulqa_few_shot each contain the per-task language-wise correlations computed from the Elo scores along with p-values, confidence intervals, the number of models for which Elo scores were available, and the number of model ranks distinguishable with about 95% confidence. Tables LABEL:tab:lmsys_pearson_alltasks_few_shot and LABEL:tab:lmsys_pearson_alltasks_few_shot contains the same information for the average of the accuracy across all tasks, and LABEL:tab:lmsys_pearson_reasoning_few_shot for the average of the accuracy across all tasks except EU21-TQA.
@@ Appendix D Influence of translation service
In this Section, we compare the OKAPI dataset with our dataset. We use the OKAPI datasets from the LM Eval Harness, which has been extended by seven languages compared to the original publication Lai et al. (2023).
D.1 Dataset overview
Table 29 compares the number of translated samples in the 11 languages that overlap with both datasets. Our dataset aims to cover all samples in different languages. For the validation splits, we cover all samples across all languages except for Hellaswag (99.93% translated). For the MMLU and ARC benchmark, the validation split is used for in-context examples in the LM Eval Harness. For the test split, we also covered all samples. The OKAPI translations cover, on average, 92.23% of the samples for all languages regardless of the split, excluding the MMLU validation split. For the MMLU validation split, 276 samples were translated on average.
Table 29: Average sample number for each task of eleven translations.
Split Original EU21 OKAPI
MMLU Val. 1531 1531 276.18 ±$\pm 8$
Test 14042 14042 13198.09 ±$\pm 101$
Hellaswag Val 10042 10035.18 ±$\pm 19$ 9276.63 ±$\pm 112$
ARC Val. 299 299 272.66 ±$\pm 0.5$
Test 1172 1172 1071 ±$\pm 2$
TruthfulQA Val. 817 817 782.09 ±$\pm 6$
D.2 Accuracy Plots
In Figures 10 , 11, and 12, we compare the accuracy of 19 LLMs on the EU21-ARC, EU21-Hellaswag, and EU21-MMLU with the OKAPI benchmarks, respectively. All models show a higher accuracy on our benchmarks regardless of the benchmark, with the only exception being the results of bloom-7b1 for the MMLU task (cf. Figure 12).
@@ D.2 Accuracy Plots
D.3 Correlations
We have calculated the correlation using the Pearson correlation coefficient for a more detailed analysis of the results from Figures 10 , 11, and 12. In Table 30, 31, and 32 we show the model performance correlation matrix clustered by model name, language, and task, respectively.
Table 30: Model Performance Correlation Matrix
(PC: Pearson Correlation Coefficient)
Model Name PC
Aya-23-8B 0.571
Meta-Llama-3-8B-Instruct 0.562
Mistral-7B-Instruct-v0.3 0.562
Meta-Llama-3-8B 0.556
gemma-7b 0.555
Llama-2-7b-chat-hf 0.553
Meta-Llama-3.1-8B 0.552
gemma-1.1-7b-it 0.552
Meta-Llama-3.1-8B-Instruct 0.549
occiglot-7b-eu5 0.548
occiglot-7b-eu5-instruct 0.546
Mistral-7B-v0.1 0.542
Mistral-7B-Instruct-v0.2 0.539
Mistral-7B-v0.3 0.536
Phi-3-mini-4k-instruct 0.534
Mistral-7B-Instruct-v0.1 0.528
Llama-2-7b-hf 0.520
bloomz-7b1 0.439
bloom-7b1 0.401
Table 31: Language Correlation Coefficients
(PC: Pearson Correlation Coefficient)
Language PC
es 0.589991
fr 0.586389
it 0.567793
sv 0.564892
da 0.557272
nl 0.555682
ro 0.546222
de 0.545930
sk 0.473632
hu 0.444939
Table 32: Task Correlation Coefficients
(PC: Pearson Correlation Coefficient)
Task PC
ARC 0.610
HellaSwag 0.543
MMLU 0.538
D.4 Quality assurance
In addition to the performance and correlation of the performance of the models, we also analyzed the translation quality using the COMET metric. Here, we performed the translations from our dataset and the OKAPI dataset for all available comparable languages. The metric we used was the COMET-KIWI metric, which is calculated using the “Unbabel/wmt22-cometkiwi-da” model [^22] https://huggingface.co/Unbabel/wmt22-cometkiwi-da. This model is reference-free and built on top of InfoXLM Rei et al. (2023). For Hellaswag and MMLU, the results of our translations are slightly higher than those of OKAPI with no evidence of significance.
Table 33: Comparison of EU21 and Okapi scores for ARC and HellaSwag benchmarks (Higher values in bold)
ARC Hellaswag MMLU
Language EU21 OKAPI EU21 OKAPI EU21 OKAPI
@@ D.4 Quality assurance
D.5 Translation examples
The following are selected examples from the EU21 and OKAPI translations of the datasets.
Table 34: Hellaswag translation examples compared between EU21 and OKAPI
@@ D.5 Translation examples
German/EU21 COMET: 0.6329 Wie man einen guten Wellensittich auswählt. Suchen Sie nach Wellensittichen, die von einem seriösen Verkäufer verkauft werden. Sie können Wellensittiche im Internet, in Kleinanzeigen oder in Zoohandlungen finden, die zum Verkauf stehen. Sie können auch einen Wellensittich aus Ihrem örtlichen Tierheim bekommen. - Die meisten Tierheime verkaufen Wellensittiche gegen eine Pauschalgebühr, so dass sich Ihre Bankkosten in Grenzen halten. Vielleicht können Sie auch die Preise vergleichen, um ein gutes Angebot zu bekommen. - Sie müssen einen angemessenen Preis finden. Im Allgemeinen sind 20-50 % des Schätzpreises angemessen. - Wo auch immer Sie Ihren Wellensittich kaufen, stellen Sie sicher, dass der Verkäufer gesunde und artgerecht gehaltene Vögel hat. Schauen Sie sich die Online-Bewertungen der Verkäufer an, die Sie finden. - Grundsätzlich müssen Sie einen guten Verkäufer finden, der bereits Vorbesitzer hatte. Erkundigen Sie sich, ob der Wellensittich den Eltern gehört hat.
@@
French/EU21 COMET: 0.5770 Une page de titre verte apparaît, avec des mots indiquant que la vidéo sera celle d’un joueur de saxophone ténor, ainsi que le site web permettant de recevoir des leçons. Un homme noir apparaît ensuite, avec un pull vert et un chapeau noir, jouant du saxophone. l’homme - bluffe à la batterie et la vidéo continue, à différents moments, à montrer des personnes jouant de l’instrument. - - continue à jouer et applaudit, serre la main de l’homme, lui lance une serviette, écrit sur la serviette et lui donne une carte. pose ensuite l’instrument sur ses genoux et commence à parler de ce qu’il fait. - saute dans une vidéo en noir et blanc et s’arrête temporairement de jouer en disant " big calibers " (gros calibres).
@@
Spanish/EU21 COMET: 0.7803 Un hombre bebe de una lata junto a un agujero en la nieve. él - nada lentamente por una masa de agua poco profunda. - habla y gesticula a la cámara. - un joven le muerde los dientes. - se acerca con pico y pala.
@@
Italian/EU21 COMET: 0.8592 Come alleviare un virus intestinale. Evitate i cibi solidi per qualche ora per permettere allo stomaco di calmarsi. Quando si avvertono i primi sintomi di un virus intestinale, è necessario lasciare che lo stomaco si stabilizzi. Mangiare cibo subito potrebbe irritare maggiormente la mucosa gastrica, peggiorando i sintomi. - Per qualche ora, evitate i cibi solidi. Bevete piccoli sorsi d’acqua o masticate del ghiaccio. - Evitare di mangiare qualcosa per almeno 10 minuti. Gli studi non hanno riscontrato che mangiare per qualche ora possa alleviare i sintomi dello stomaco. - Cercare di mangiare un pasto solido qualche ora prima dell’attacco di nausea non farà altro che peggiorare i sintomi. Se vi sentite male dopo l’inizio dei sintomi, smettete di mangiare. - Cercate di mangiare qualche ora prima del lavoro o della scuola, soprattutto se vi sentite male. Anche muoversi un po’ può ridurre la nausea e aiutare lo stomaco a stabilizzarsi.
@@
Swedish/EU21 COMET: 0.8174 En man med hörlurar befinner sig i en studio och talar i en mikrofon. Andra män visas, varav en börjar spela munspel. han - slutar spela och säger något som får dj:n att skratta. - ses leka intensivt med kameran. - leker med en jack o lantern och lämnar rummet. - verkar mycket nöjd med den information han har samlat in när han låter programledaren tala.
@@
Hungarian/EU21 COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
@@
Slovak/EU21 COMET: 0.8336 Muž beží a pritom drží tyč na dráhe. on - používa tyč na skok vo vzduchu. - beží po dráhe a obopína nohy okolo tyče. - vbehne do pieskoviska a začne skákať. - dopadne na hromadu piesku.
@@
Danish/EU21 COMET: 0.8450 En lille pige løber væk fra pinataen, mens hun holder pinden. Barnet vender tilbage og rammer pinataen én gang. personerne - fortsætter med at fjerne nøgleingredienser fra poserne. - jubler over barnet, der bliver ramt. - ser til, mens pigen kaster pinden to gange. - alle klapper, og pigen smiler og vender sig mod kameraet.
@@
Hungarian/EU21 COMET: 0.8367 Hogyan kell egy nem felekezeti temetési szertartást levezetni. Beszélj a családdal. Hívja fel a családot, és egyeztessen velük egy találkozót, lehetőleg az otthonukban. A nyilvános találkozás kötöttségek esetén rendben van, de ha ez a megállapodás, próbáljon meg nem két-három embernél többel találkozni. - Ha a szertartás templomban lesz, ne fáradjon azzal, hogy kint üljön egy parkban lévő padon. Ha lelkész vagy bába, a legjobb, ha a szokásos istentiszteleti időre érkezik. - Természetesen csak egy családtaggal és az egyik gyermekükkel együtt nem baj. Ha a temetéseden senki sem hajlandó részt venni, az sem baj. - Az emberek nem mindig örülnek a szertartásvezető apró cselekedeteinek. Egyes esetekben a család részeggé, izgatottá vagy erőszakossá válhat a szertartás alatt vagy körül. - Ha egy temetkezési vállalat gondoskodik egyéb előkészületekről, a temetkezési vállalatnak lehet, hogy van olyan hely a ravatalozóban, ahol találkozhatnátok. Hívja meg a családot, hogy más családtagok és közeli barátok is jelen legyenek, ha kívánják, de legyen tekintettel az ő igényeikre és a térbeli megfontolásokra.
@@
Table 34: Hellaswag translation examples compared between EU21 and OKAPI
Table 35: ARC and MMLU translation examples each compared between EU21 and OKAPI
@@
German/EU21 COMET: 0.8345 Auf wen trafen die anatomisch modernen Menschen bei ihrer ersten Ankunft im Nahen Osten? - Neandertaler, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo erectus, die evolutionären Nachfahren der vormodernen menschlichen Bewohner Europas und Asiens - Homo sapiens, der sich unabhängig von den Menschen in Asien entwickelt hatte und dort vor 200.000 Jahren ankam - Alle oben genannten
@@
German/EU21 COMET: 0.8546 Eine Klasse plant eine Untersuchung, um herauszufinden, welche Marke von Glühbirnen am längsten hält. Welcher dieser Schritte sollte zuerst erfolgen? - Wiederhole die Untersuchung. - Schreibe einen Bericht über die Ergebnisse. - Fertige eine Tabelle an, um die Daten festzuhalten. - Beobachte täglich die Glühbirnen.
@@
Table 35: one random ARC and MMLU translation examples each compared between EU21 and OKAPI
Appendix E Other
E.1 License
MMLU, Hellaswag and GSM8K were released under MIT license; ARC under CC-BY-SA-4.0 license; TruthfulQA under Apache 2.0 license.
EU20-MMLU, EU20-HellaSwag, EU20-GSM8K and EU20-TruthfulQA will be released with an permissive license, we ensure that all original licenses and notices are included. EU20-ARC will be released under the CC-BY-SA-4.0 license under the exactly the same terms as the original dataset ARC.
E.2 Use of AI Assistants
We acknowledge the use of ChatGPT-4 to generate code snippets for the Figures 1-3. The code was then extended and carefully reviewed.